Consider the titration of a 24.5 mL sample of 0.115 M RbOH with 0.110 M HCL . Determine each of the following:

A) initial pH

B) the volume of added acid required to reach the equivalence point

C) the pH at 5.1 mL of added acid

D) the pH at the equivalence point

E) the pH after adding 5.6 mL of acid beyond the equivalence point
Answer #1

A)when 0.0 mL of HCl is added

Given:

M(HCl) = 0.11 M

V(HCl) = 0 mL

M(RbOH) = 0.115 M

V(RbOH) = 24.5 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.11 M * 0 mL = 0 mmol

mol(RbOH) = M(RbOH) * V(RbOH)

mol(RbOH) = 0.115 M * 24.5 mL = 2.8175 mmol

We have:

mol(HCl) = 0 mmol

mol(RbOH) = 2.818 mmol

0 mmol of both will react

remaining mol of RbOH = 2.818 mmol

Total volume = 24.5 mL

[OH-]= mol of base remaining / volume

[OH-] = 2.818 mmol/24.5 mL

= 0.115 M

use:

pOH = -log [OH-]

= -log (0.115)

= 0.9393

use:

PH = 14 - pOH

= 14 - 0.9393

= 13.0607

Answer: 13.06

B)

M(RbOH)=0.115 M

M(HCl)=0.11 M

V(RbOH)=24.5 mL

According to balanced reaction:

1*number of mol of RbOH =1*number of mol of HCl

1*M(RbOH)*V(RbOH) =1*M(HCl)*V(HCl)

1*0.115 M *24.5 mL = 1*0.11M *V(HCl)

V(HCl) = 25.6 mL

Answer: 25.6 mL

C)when 5.1 mL of HCl is added

Given:

M(HCl) = 0.11 M

V(HCl) = 5.1 mL

M(RbOH) = 0.115 M

V(RbOH) = 24.5 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.11 M * 5.1 mL = 0.561 mmol

mol(RbOH) = M(RbOH) * V(RbOH)

mol(RbOH) = 0.115 M * 24.5 mL = 2.8175 mmol

We have:

mol(HCl) = 0.561 mmol

mol(RbOH) = 2.818 mmol

0.561 mmol of both will react

remaining mol of RbOH = 2.257 mmol

Total volume = 29.6 mL

[OH-]= mol of base remaining / volume

[OH-] = 2.257 mmol/29.6 mL

= 7.623*10^-2 M

use:

pOH = -log [OH-]

= -log (7.623*10^-2)

= 1.1179

use:

PH = 14 - pOH

= 14 - 1.1179

= 12.8821

Answer: 12.88

D)

This is titration of strong acid and strong base.

pH at equivalence point will be 7.00 as the solution would be neutral

Answer: 7.00

E)

1)when 31.2 mL of HCl is added

Given:

M(HCl) = 0.11 M

V(HCl) = 31.2 mL

M(RbOH) = 0.115 M

V(RbOH) = 24.5 mL

mol(HCl) = M(HCl) * V(HCl)

mol(HCl) = 0.11 M * 31.2 mL = 3.432 mmol

mol(RbOH) = M(RbOH) * V(RbOH)

mol(RbOH) = 0.115 M * 24.5 mL = 2.8175 mmol

We have:

mol(HCl) = 3.432 mmol

mol(RbOH) = 2.818 mmol

2.818 mmol of both will react

remaining mol of HCl = 0.6145 mmol

Total volume = 55.7 mL

[H+]= mol of acid remaining / volume

[H+] = 0.6145 mmol/55.7 mL

= 1.103*10^-2 M

use:

pH = -log [H+]

= -log (1.103*10^-2)

= 1.9573

Answer: 1.96